OpenStudy (anonymous):

sherry spent 4 hours excercising last week. if sherry spent 5 over 6 of the time she spenmt jogging how much time did she spent jogging?help please i dont get and if you cant give the answer please expolain.

OpenStudy (anonymous):

You question is typed wrong. Retype it and I can help.

OpenStudy (anonymous):

sherry spent 4 hours exercising last week.if 5 (over) 6 <--(fraction) of the time she spent jogging, how much time did she spend joging.write answer in simplest form.

OpenStudy (anonymous):

200 minutes or 3 hours 20 minutes. 5/6 = 50/60 = 50 minutes of every hour times 4 hours = 200 minutes, divide by 60 = 3 hours 20 minutes